A proofreader is interested in finding the probability that the number of mistakes in a page will be less than 10. From his past experience he finds that out of 3600 pages he has proofed, 200 pages contained no errors, 1200 pages contained 5 errors, and 2200 pages contained 11 or more errors. Can you help him in finding the required probability? 


1
Expert's answer
2022-04-29T10:21:22-0400

Let "X=" the number of mistakes in a page.


"P(X<10)=\dfrac{200+1200}{3600}=\dfrac{7}{18}"